### v1

# Revela: Dense Retriever Learning via Language Modeling

Abstract
Dense retrievers play a vital role in accessing external and specialized knowledge to augment language models (LMs). Training dense retrievers typically requires annotated query-document pairs, which are costly and hard to obtain in special domains, like code—motivating growing interest in self-supervised retriever learning. Since LMs are trained to capture token-level dependencies through a self-supervised learning objective (i.e., next token prediction), we can analogously cast retrieval as learning dependencies among chunks of tokens. This analogy naturally leads to the question: How can we adapt self-supervised learning objectives in the spirit of language modeling to train retrievers?
To answer this question, we introduce Revela, a unified and scalable training framework for self-supervised retriever learning via language modeling. Revela models semantic dependencies among documents by conditioning next token prediction on both local and cross-document context through an in-batch attention mechanism. This attention is weighted by retriever-computed similarity scores, enabling the retriever to be optimized as part of language modeling. We evaluate Revela on both general-domain (BEIR) and domain-specific (CoIR) benchmarks across various retriever backbones. At a comparable parameter scale, Revela outperforms the previous best method with absolute improvements of 5.2% (18.3% relatively) and 5.6% (14.4% relatively) on NDCG@10, respectively, underscoring its effectiveness. Performance increases with model size, highlighting both the scalability of our approach and its promise for self-supervised retriever learning.
[…]
1 Introduction
Central to information retrieval are dense retrievers (reimers2019sentence, ; karpukhin-etal-2020-dense, ; ma2024fine, ), which map queries and documents into high-dimensional vector spaces and determine relevance through similarity calculations. Typically, these models rely on carefully annotated query-document pairs and hard negatives for training. However, creating such high-quality training data requires substantial human annotation, which is labor-intensive and difficult to scale in real-world, domain-specific scenarios such as law feng-etal-2024-legal and programming jimenez2024swebench . This limitation stimulates the interest within the community to explore self-supervised approaches for training retrievers directly from unannotated raw texts, i.e., self-supervised retriever learning DBLP:journals/tmlr/IzacardCHRBJG22 .
Modern LMs (grattafiori2024llama, ), a successful case of self-supervised learning, are typically pretrained with the next-token prediction (NTP) paradigm, modeling dependencies among tokens within a single sequence. Analogously, retriever aims to model relationships among larger units—chunks of tokens—capturing more macroscopic dependencies. This motivates us to raise a key question: How can we learn a retriever within a self-supervised learning framework of language modeling? While NTP implicitly identifies the most relevant parts in the context during generation, conditioning language modeling of one sequence on others can serve as an effective proxy for modeling inter-sequence relationships. This can offer a novel and principled approach to self-supervised retriever learning.
In this work, we introduce a self-supervised retriever learning paradigm – Dense Retriever Learning via Language Modeling, abbreviated as Revela.  [^1] Revela merges re- (“retriever”) with vela (“sails” in Latin) to evoke a self-supervised method where retrievers rig the “sails” of language modeling to capture the ambient wind of raw text for label-free learning. As illustrated in Fig.1, Revela trains retrievers by simultaneously optimizing retrievers and LMs; different from conventional NTP, Revela learns the probability of a token given both the prefix in this sequence and all other sequences in the batch. Specifically, in addition to classical self-attention which restricts NTP to individual sequences, an in-batch attention mechanism enables sequences to attend to their neighbors in the same batch during language modeling. In this process, the retriever provides inter-sequence dependencies that modulate the in-batch attention weights, allowing it to be optimized with the LM during training. We split raw texts into chunks within each document and put these chunks into the same batch, motivated by the idea of hard negative samples in contrastive learning (xiong2020approximatenearestneighbornegative, ).
Figure 1: The framework of Revela. Similarity scores between chunks within a batch computed by dense retrievers are used as attention weights for in-batch attention inside Revela’s transformer block. The retriever is trained by optimizing the language modeling objective, i.e., NTP. For instance, when modeling the red sequence, the similarity to the purple chunk will be increased, with related segments highlighted in bold, underline, and italic.
To assess the effectiveness of Revela across diverse retrieval scenarios, we evaluate it on two complementary benchmarks: BEIR (thakur2021beir, ), which covers general-domain retrieval tasks, and CoIR (li2024coir, ), a comprehensive benchmark tailored for code retrieval. Specifically, Revela is trained on Wikipedia for general retrieval and on a set of code-related corpora for code retrieval wang2024coderagbench , using pretrained transformers ranging from 135M to 1B parameters as retriever backbones, with an LM for language modeling. We observe that, with a comparable number of parameters, Revela surpasses REPLUG (shi-etal-2024-replug, ) by an absolute margin of 5.3% on BEIR and 5.4% on CoIR in terms of NDCG@10, corresponding to relative improvements of 18.3% and 14.3%, respectively. These results demonstrate the effectiveness of Revela as a self-supervised retriever learning solution. We also observe that scaling Revela results on the size of retriever backbones and LMs in higher performance gain over baselines. Moreover, Revela maintains consistent performance on both BEIR and CoIR when trained on a mixed-domain dataset. These results collectively highlight Revela’s promising scalability and generalization across both model size and data scale.
To this end, we summarize our contributions as follows:
  • We introduce Revela, a self-supervised framework that trains a retriever via language modeling using an additional in-batch attention mechanism, where next-token prediction is conditioned on both the input sequence and others within the same batch.
  • On the BEIR and CoIR benchmarks, Revela surpasses REPLUG, achieving relative NDCG@10 improvements of 18.3% and 14.3%, respectively, underscoring its effectiveness.
  • Revela benefits from both parameter and data scaling: larger retriever and LM backbones yield additional gains, while consistent performance across two benchmarks with mixed-domain training data highlights its generalization.
2 Related Works
Self-supervised Retriever Learning
Dense retrievers are typically trained with query-document pairs, requiring extensive human annotation. Given the abundance of unlabeled corpora, a key challenge in the community is: How can we train a dense retriever in a self-supervised manner?
One line of works leverages weak supervision within documents. Contriever (DBLP:journals/tmlr/IzacardCHRBJG22, ) employs contrastive learning, treating passages from the same document as positives and using in-batch negatives, while Co-Condenser (gao-callan-2022-unsupervised, ) warms up Condenser gao-callan-2021-condenser with Wikipedia hyperlinks as pseudo labels. However, these methods are limited by overly simplistic assumptions that structure-associated passages are labeled as pseudo-positives, failing to capture complex relational patterns. Another common strategy is to distill from other retrievers, exemplified by LaPraDoR xu-etal-2022-laprador , which combines dense retrieval with BM25 robertson2009probabilistic through Lexicon-Enhanced Dense Retrieval. Last but not least, autoencoding methods, such as TSDAE (wang-etal-2021-tsdae-using, ) and RetroMAE (xiao-etal-2022-retromae, ), learn sentence embeddings by reconstructing sentences from their neural representations. Autoencoding often lacks pairwise or contrastive supervision, which can result in overfitting to low-level details (steck2020autoencoders, ). Inspired by the aforementioned approaches, our method trains retrievers by modeling dependencies within a document, generating each sequence conditioned on others within the same batch.
LM-guided Retriever Learning
LMs (e.g., GPT (NEURIPS2020_1457c0d6, ), LLaMA (touvron2023llama, )) are trained on trillions of tokens with billions of parameters, enabling strong generalization and improving retriever learning through richer semantics and training signals.
LMs can facilitate query and document expansion, as demonstrated by approaches like Query2Doc (wang-etal-2023-query2doc, ) and Doc2Query (gospodinov2023doc2query, ). However, leveraging such capabilities often requires advanced models, leading to high computational costs. Augmenting LMs with relevant information from external knowledge stores not only improves performance in various NLP tasks but also enhances retriever learning. Atlas (izacard2023atlas, ) utilizes cross-attention scores between retrieved documents and the generated output as signal to train the retriever. However, it is based on an encoder-decoder architecture, which diverges from the prevailing trend of decoder-only LMs (touvron2023llama, ; achiam2023gpt, ). Additionally, it requires periodic index updates (reindexing) (guu2020retrieval, ), leading to significant computational costs. In contrast, our work focuses on leveraging the widely adopted decoder-only architecture while eliminating the need for reindexing by modeling relationships between chunks within the same document.
With the rise of decoder-only LMs, REPLUG (shi-etal-2024-replug, ) enhances retrieval by prepending retrieved documents to the queries, training retrievers to produce query-document similarity aligned with the LM’s perplexity. However, the perplexity of frozen LMs is often poorly calibrated (geng-etal-2024-survey, ), resulting in suboptimal retriever learning. This issue can be optimized in Revela where retrievers and LMs are updated jointly during language modeling.
Domain-specific Retrieval
In pre-training corpora, domain-specific knowledge is both scarce and rapidly evolving (grossmann2023ai, ; wen2025synthetic, ), making effective retrieval in specialized domains critically challenging. To enhance the adaptability of dense retrievers across domains, researchers have explored continual learning (sachan-etal-2021-end, ; oguz-etal-2022-domain, ) and task-aware training (cheng-etal-2023-task, ). However, these approaches still rely on query-document pairs from domain-specific datasets. Another approach seeks to simplify domain-specific retrieval for general-purpose dense retrievers. Cai et al. cai-etal-2024-mixgr propose a divide-and-conquer strategy through a mixed-granularity retrieval framework, significantly enhancing dense retriever performance in scientific domains. Our work demonstrates Revela’s domain adaptation capability through language modeling on domain-specific raw texts.
3 Revela: Dense Retriever Learning via Language Modeling
3.1 Training Objectives
LM training typically includes the maximization of NTP for token sequences. Given a batch of documents $\{D_{1},D_{2},\dots,D_{B}\}$ and an LM parameterized by $\Phi$, classical NTP on the token $x_{l}^{i}$ in $D_{i}=\{x_{1}^{i},\dots,x_{L}^{i}\}$, can be calculated as, where $x_{<l}^{i}$ denotes the tokens preceding $x_{l}^{i}$ in $D_{i}$
$P(x_{l}^{i})=P_{\Phi}(x_{l}^{i}\mid x_{<l}^{i}).$ (1)
Figure 2: Revela’s architecture. With a specific attention map, the embeddings of in-batch attention $\{\mathrm{h}_{i}^{l}\}_{i=1}^{B}$ can attend to the self-attention $\{\mathrm{e}_{i}^{l}\}_{i=1}^{B}$.
As show in Fig. 2, in Revela, the NTP for sequence $i$ is conditioned not only on its own preceding context $x_{<l}^{i}$ but also on every other document in the batch, $\{D_{j}\}_{j\neq i}$. Specifically, we introduce a new attention mechanism in the transformer blocks, i.e., in-batch attention (Sec. 3.2), which is supported by a specific mask design (Sec. 3.4). The attentions are weighted by the inter-sequence similarity computed by the retriever parameterized by $\Theta$, i.e., $\mathrm{Sim}(D_{i},D_{j})$ (Sec. 3.3), where $\sum_{j\neq i}^{B}\mathrm{Sim}(D_{i},D_{j})=1$. The retriever is dynamically optimized as the similarity is updated jointly with the NTP objective
$P_{R}(x_{l}^{i})=P_{\Phi,\Theta}\!\bigl{(}x_{l}^{i}\mid x_{<l}^{i},\{D_{j}\}_{j\neq i}\bigr{)}$ (2)
3.2 In-batch Attention
Revela augments the transformer block with an in-batch attention mechanism to incorporate context from other sequences. We denote the output of $D_{i}$ at the $l$-th layer as $[\mathrm{e}_{i}^{l};\mathrm{h}_{i}^{l}]$, where $\mathrm{e}_{i}^{l},\mathrm{h}_{i}^{l}\in\mathbb{R}^{L\times d}$ represent the outputs from self-attention and in-batch attention, respectively. For simplicity, we omit layer normalization and feed-forward layers, so the input to the two attention modules in the $l$-th layer is $[\mathrm{e}_{i}^{l-1};\mathrm{h}_{i}^{l-1}]$.
Standard Self-Attention
In the $l$-th layer of the blocks, the self-attention mechanism computes
$Q_{i}^{e}=\mathrm{e}^{l-1}_{i}W^{Q},\quad K_{i}^{e}=\mathrm{e}^{l-1}_{i}W^{K},\quad V_{i}^{e}=\mathrm{e}^{l-1}_{i}W^{V},$ (3)
where $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d\times d}$ are learnable projection matrices. For multi-head attention with $H$ heads (each of dimension $d_{H}=d/H$), the standard causal attention is computed as
$\mathrm{e}_{i}^{l}=\text{softmax}\!\left(\frac{Q_{i}^{e}K_{i}^{e\top}}{\sqrt{d_{H}}}\right)V_{i}^{e}.$ (4)
In-batch attention
combines standard self-attention with cross-document attention. The embeddings of $k$-th token in $D_{i}$ is obtained by (1) the prefix of $D_{i}$, and (2) the other documents $\{D_{j}\}_{j\neq i}$ based on their similarity to $D_{i}$. This encourages $D_{i}$ to selectively attend to more relevant documents based on learned retrieval signals. For the contribution from the prefix of $D_{i}$, the self-attention output $\mathrm{s}_{i}^{l}$ in the in-batch attention uses the same projection matrices as above
$Q_{i}^{h}=\mathrm{h}_{i}^{l-1}W^{Q},\quad K_{i}^{h}=\mathrm{h}_{i}^{l-1}W^{K},\quad V_{i}^{h}=\mathrm{h}_{i}^{l-1}W^{V},\quad\mathrm{s}_{i}^{l}=\text{softmax}\!\left(\frac{Q_{i}^{h}K_{i}^{h\top}}{\sqrt{d_{H}}}\right)V_{i}^{h}.$
Cross-document attention enables $D_{i}$ to attend to other documents $D_{j}$ using cached keys $K_{j}^{e}$ and values $V_{j}^{e}$. With a full attention mask, the output $\mathrm{b}^{l}_{ij}$ is computed as
$\mathrm{b}^{l}_{ij}=\text{softmax}\!\left(\frac{Q_{i}^{h}K_{j}^{e\top}}{\sqrt{d_{H}}}\right)V_{j}^{e}.$ (5)
Normalization by value norms prevents any single token from dominating the attention by scaling the output $\mathrm{b}^{l}_{ij}$ with the norm of $V_{j}^{e}$, thereby encouraging the cross-document attention to focus on sequence-level semantics izacard2023atlas . This operation, referred to as V-normalization, is computed as
$\tilde{\mathrm{b}}^{l}_{ij}=\frac{\mathrm{b}^{l}_{ij}}{N_{ij}+\epsilon},\quad\text{where }N_{ij}=\text{softmax}\!\left(\frac{Q_{i}^{h}K_{j}^{e\top}}{\sqrt{d_{H}}}\right)\|V_{j}^{e}\|_{2}.$ (6)
where $\epsilon$ is a small constant for numerical stability.
Weighting by cross-document similarity aggregates the normalized attention outputs $\tilde{\mathrm{b}}^{l}_{ij}$ using the similarity scores $\mathrm{Sim}(D_{i},D_{j})$ computed by the retriever
$\mathrm{b}^{l}_{i}=\sum_{\begin{subarray}{c}j=1,j\neq i\end{subarray}}^{B}\mathrm{Sim}(D_{i},D_{j})\,\tilde{\mathrm{b}}^{l}_{ij}.$ (7)
Combined output integrates the results of self-attention and cross-document attention to form the final output of the in-batch attention
$\mathrm{h}^{l}_{i}=\mathrm{s}^{l}_{i}+\mathrm{b}^{l}_{i}$ (8)
3.3 Similarity Computation
Given a batch of $B$ documents $\{D_{i}\}_{i=1}^{B}$, the retrieval mechanism proceeds in three steps. First, each document $D_{i}$ is encoded into an embedding $\mathbf{h}_{i}\in\mathbb{R}^{d_{E}}$ using an encoder $E_{\Theta}$, such that $\mathbf{h}_{i}=E_{\Theta}(D_{i})$. Second, the embeddings are normalized and pairwise cosine similarities are computed: $\tilde{\mathbf{h}}_{i}=\mathbf{h}_{i}/\|\mathbf{h}_{i}\|_{2}$, and the similarity score between documents $i$ and $j$ is given by $S_{ij}=\tilde{\mathbf{h}}_{i}^{\top}\tilde{\mathbf{h}}_{j}$. Third, temperature-scaled softmax with temperature $\tau$ is applied to obtain probabilities across documents
$\mathrm{Sim}(D_{i},D_{j})=\frac{\exp(S_{ij}/\tau)}{\sum_{k\neq i}\exp(S_{ik}/\tau)}.$
The resulting weights $\mathrm{Sim}(D_{i},D_{j})\in\mathbb{R}^{B\times B}$ capture cross-document similarities and are used to condition language modeling on relevant in-batch documents as presented above.
3.4 Implementation Details
As described earlier, Revela adapts the classical transformers by additionally including in-batch attention, which builds upon standard self-attention, as shown in Eq. 5. For the minimum modifications to the existing transformer’s implementation, we take the computation of $\mathrm{e}$ and $\mathrm{h}$ as duplicating documents and adjusting the attention mask, as illustrated in Fig. 2. In this way, the embeddings $\{\mathrm{h}_{i}^{l}\}_{i=1}^{B}$ produced by in-batch attention can be obtained by applying full attention over the self-attention outputs $\{\mathrm{e}_{i}^{l}\}_{i=1}^{B}$ and aggregating them.
4 Experiments
4.1 Experimental Setups
Evaluation Benchmarks
To comprehensively evaluate our proposed framework, we benchmark Revela on two diverse datasets: BEIR (thakur2021beir, ), a heterogeneous benchmark covering multiple domains for general information retrieval, and CoIR (li2024coir, ), a comprehensive benchmark designed for code-specific retrieval tasks. A more detailed introduction is listed in Appendix A.1.
Training Data
We use two different recipes of raw texts for the general and code-related retrieval mentioned above. The examples are presented in Appendix A.2. In this section, where we study the data independently, we also study the impact of the mixture of these two datasets in Sec. 5.2.
  • BEIR: We segment the passages in the Wikipedia corpus [^2] https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus into chunks of at most 120 words, each comprising complete sentences. Given a set of passages $\{d_{1},d_{2},\ldots,d_{n}\}$, where each passage $d_{i}$ is divided into a sequence of chunks $(d_{i1},d_{i2},\ldots,d_{im_{i}})$, the chunks are interleaved in the order $(d_{11},d_{21},\ldots,d_{n1},d_{12},d_{22},\ldots)$, and grouped sequentially into batches of size 16. In total, we sample 320,000 batches for training.
  • CoIR: Similarly, we chunk a set of code-related corpora (wang2024coderagbench, ), including the posts in the StackOverflow forum (weber2024redpajama, ), online tutorials (overwijk2022clueweb22, ), and library documentations (zhoudocprompting, ), for CoIR. The batch size is maintained as 16. Overall, there are 358,763 training batches.
Models
In Revela, we jointly train a retriever and an LM using self-supervised learning of LM, i.e., NTP. We adopt LLaMA-3.2-1B (grattafiori2024llama, ) as the LM. To ensure a fair comparison across diverse unsupervised baselines and retriever architectures, we adopt a range of LMs with parameter sizes from 0.1B to 1B as the backbone models for the retrievers: SmolLM2-135M (allal2025smollm2, ), Qwen2.5-0.5B (yang2024qwen2, ), and LLaMA-3.2-1B. We follow the approach used in RepLLaMA (ma2024fine, ). We append the <eos> token to each sentence and use its corresponding embedding as the sentence representation. Additionally, we prepend the prefixes "Query: " and "Passage: " to queries and passages, respectively. For more details about model checkpoints, please refer to Appendix A.4.
Baselines
We include representative self-supervised baselines: Contriever (DBLP:journals/tmlr/IzacardCHRBJG22, ) trains a BERT-based retriever via contrastive learning on unsupervised pairs; RetroMAE (xiao-etal-2022-retromae, ) reconstructs masked passages from compressed representations; LaPraDor (xu-etal-2022-laprador, ) combines latent-pair contrastive pretraining on C4 (JMLR:v21:20-074, ) with BM25-enhanced dense retrieval. REPLUG (shi-etal-2024-replug, ) distills supervision signals from frozen language models into retrievers, i.e., modeling chunk-chunk similarity within a batch by distilling the perplexity from an LM, using one chunk as prior context for the other. We select REPLUG [^3] For more details about REPLUG replication, please refer to Appendix A.3. as the main baseline for three reasons: (1) REPLUG consists of a retriever and a decoder-only LM, which is similar to Revela. Other approaches adopt different architectures—for instance, the encoder-decoder structure in Atlas (izacard2023atlas, )—which may divert attention from our primary focus on the joint training of retrievers and LMs. (2) Architectural similarity ensures that comparisons remain generalizable across different scales. (2) As reported by Shi et al. shi-etal-2024-replug , REPLUG outperforms most prior methods and can serve as a representative of previous research in this line of work.
When benchmarking CoIR, we also include several supervised retrievers, as well as API-based models such as OpenAI-Ada-002 (Ada-2) and Voyage-Code-002 (Voyage-2), following original setups (li2024coir, ). Supervised retrievers include GraphCodeBERT (GCBERT) (guographcodebert, ) and UniXcoder (guo-etal-2022-unixcoder, ), which are finetuned on code-related datasets. For more details about the Huggingface checkpoints, please refer to Appendix A.4.
Experimental Details
For both retrievers and LMs, we apply LoRA with a rank of 256. Training uses the temperature $\tau$ $1\mathrm{e}{-4}$, the small constant $\epsilon$ $1\mathrm{e}{-6}$, a learning rate of $1\mathrm{e}{-4}$, and 100 warmup steps, following the WarmupDecayLR schedule. We train on 4× A100 80GB GPUs with a gradient accumulation step size of 8. Passages are truncated to 160 tokens, and bf16 mixed precision is enabled. We finetune the models for one epoch, therefore, there are 10,000 steps on Wikipedia and around 11,000 steps on code-related texts.
Specifically, when calculating the similarity between chunks derived from Wikipedia, we follow REPLUG—using only the first half of the chunk to compute similarity with other chunks. This design is motivated by the inferential semantics inherent in natural language retrieval tasks. In contrast, for code retrieval tasks, we retain the full chunk when computing similarity, as these tasks rely more heavily on precise semantic matching compared to natural language.
Table 1: Performance of self-supervised retriever models on BEIR datasets (NDCG@10, %). Bold marks the best score per dataset for unsupervised methods. RepLLaMA (ma2024fine, ), a 1B supervised model trained on MsMarco, is reported for reference only and is not considered when highlighting the best.
Dataset BM25 Contriever RetroMAE LaPraDor REPLUG Revela REPLUG Revela REPLUG Revela Revela+ BM25 RepLLaMA
Model Size 0.1B 0.1B 0.1B 0.1B 0.1B 0.5B 0.5B 1B 1B 1B 1B
ArguAna 48.7 44.3 32.3 44.6 28.1 39.4 34.7 41.4 39.5 45.1 42.2 50.6
C-FEVER 13.6 7.2 5.9 12.2 14.4 15.6 15.3 13.5 14.1 15.6 26.5 27.7
DBPedia 29.9 27.0 4.8 25.0 13.3 18.3 18.9 21.4 19.5 27.6 42.4 42.4
FEVER 48.1 27.2 4.7 33.6 35.1 35.1 47.3 50.9 48.3 60.4 65.7 78.9
FiQA 25.1 12.4 2.8 19.8 14.4 19.1 19.6 26.6 22.0 29.6 41.0 40.1
HotpotQA 56.9 41.0 20.9 30.4 26.7 38.5 36.4 50.6 42.4 56.0 68.8 66.5
NFCorpus 32.1 27.1 6.2 30.4 20.8 23.5 25.8 26.5 26.7 27.2 30.6 35.6
NQ 28.5 18.1 3.7 18.0 18.3 21.2 26.5 29.8 27.9 33.9 44.4 57.8
Quora 80.4 83.4 56.8 78.7 72.3 80.9 76.4 83.1 82.4 83.5 86.7 84.4
SCIDOCS 15.8 10.9 4.1 13.4 9.6 12.1 13.5 14.8 14.5 16.3 27.0 17.8
SciFact 68.7 59.1 30.9 49.9 55.9 57.9 63.9 65.2 67.1 71.8 75.9 74.3
TREC-C 62.2 18.2 15.7 22.9 48.2 60.2 44.5 58.4 39.6 59.8 26.7 83.1
Touche 33.1 7.2 0.4 8.9 11.6 14.4 14.5 21.0 13.5 23.3 47.8 32.8
Mean 41.8 29.5 14.6 29.8 28.4 33.6 33.6 38.7 35.2 42.3 48.1 53.2
Table 2: Performance on code-related retrieval (NDCG@10, %). Gray indicates supervised models trained on code-specific data. Bold marks the best score among all non-proprietary models.
Dataset BM25 Contriever GCBERT UniXcoder REPLUG Revela REPLUG Revela REPLUG Revela Ada-2 Voyage-2
Model Size 0.1B 0.1B 0.1B 0.1B 0.1B 0.5B 0.5B 1B 1B
Apps 4.8 5.2 0.1 1.4 5.2 6.3 9.5 19.0 13.9 16.8 8.7 26.5
CosQA 15.6 13.2 3.6 25.1 16.5 24.2 17.7 27.6 20.1 27.7 28.9 29.8
ST2SQL 25.0 38.6 10.4 50.5 45.0 44.7 53.4 53.3 53.9 53.4 58.3 69.3
SN 40.9 33.2 3.0 60.2 35.0 43.2 42.8 52.6 50.8 60.1 74.2 81.8
SNCC 54.0 46.8 18.3 58.4 52.7 58.9 59.5 63.6 62.8 67.0 53.3 72.8
TransC 47.8 42.5 12.8 41.8 57.0 60.6 53.5 71.3 61.5 74.4 26.0 27.3
TransDL 34.4 33.6 31.7 31.0 34.0 32.9 31.2 34.1 33.3 33.7 17.7 28.7
SOQA 70.2 59.9 20.3 44.7 62.9 68.3 72.4 81.0 76.3 83.8 74.2 81.8
F-ST 68.1 44.5 6.9 36.0 51.1 65.0 59.4 73.8 66.0 74.5 47.1 65.4
F-MT 59.2 38.7 13.1 24.2 31.4 43.1 37.9 52.3 42.9 54.6 17.7 28.7
Mean 42.0 35.6 12.0 37.3 39.1 44.7 43.7 52.9 48.2 54.6 45.6 56.3
4.2 Experimental Results
Table 1 and Table 2 report the performance of Revela and baseline methods on BEIR and CoIR, highlighting Revela’s superior effectiveness. Specifically, with a comparable number of parameters (around 0.1B), Revela achieves an absolute gain of 4.1% on BEIR and 9.1% on CoIR over Contriever, one of the most popular self-supervised retriever. Moreover, Revela’s performance continues to improve as the retriever size increases to 0.5B and 1B.
Revela consistently outperforms REPLUG across models of varying sizes. On the BEIR benchmark, Revela -1B achieves an average score of 42.3 across 13 datasets, compared to 35.2 for REPLUG-1B. Similarly, on the CoIR benchmark, Revela 1B achieves 54.6, surpassing REPLUG’s 48.2. Given the large evaluation sets (33K queries for BEIR and 162K for CoIR), the improvements are statistically significant, with a z-test yielding a p-value below $10^{-9}$.
Combining the retrieval results of BM25 and Revela using Reciprocal Rank Fusion (RRF) (cormack2009reciprocal, ) effectively narrows the gap with the supervised model RepLLaMA, improving NDCG@10 by 5.8 points. On out-of-domain tasks such as SciDocs, SciFact, and Quora, this combination even surpasses RepLLaMA, highlighting the complementary strengths of BM25 and Revela. For the details about RRF implementation, please refer to Appendix A.2
In Table 2, Revela is shown to beat some supervised models across different sizes on CoIR and some proprietary models. Although Revela does not yet match the performance of Voyage-Code-002—the strongest model benchmarked on CoIR—we anticipate that scaling up the retriever backbone will further narrow the gap, and potentially surpass the closed-source models.
5 Analysis
Beyond Revela’s effectiveness shown earlier, we will explore its scalability from two aspects: LM size (Sec. 5.1) and training batch configuration (Sec. 5.2). We also validate the impact of our specific design on batch construction, batch sizes (Sec. 5.2), and model architecture (Sec. 5.3).
5.1 Larger LMs can help out-of-domain retrieval tasks
As observed in the previous section, retriever performance improves with larger backbone models. In this section, we further investigate how the size of the LM—another key component in training—affects the retriever’s effectiveness. In addition to LLaMA-3.2-1B, we also use SmolLM2-135M and Qwen2.5-0.5B as LMs, each paired with retrievers of three different sizes. All models are trained using the same experimental setup described in Sec. 4, on both training corpora.
As shown in Fig. 3, larger LMs do not show a consistent advantage in retrieval performance on BEIR–the general-domain benchmark. However, a clear positive trend is observed on the CoIR benchmark. We expect that seamlessly incorporating larger LMs will further enhance Revela’s adaptability to specific domains. For detailed results on each dataset, as well as an analysis of trainable parameters (i.e., LoRA rank) as another aspect of training size, please refer to Appendix B.1.
Figure 3: Performance comparison on BEIR and CoIR benchmarks using various combinations of retrievers and LMs. For code retrieval tasks, larger LMs consistently yield greater gains in retriever performance, whereas this trend is not observed for general-domain retrieval.
5.2 Training Dataset Configuration
Figure 4: Average performance (NDCG@10, %) on the BEIR and CoIR benchmarks across different dataset configurations (vanilla, randomly constructed, and mixed training corpus). Best results are highlighted in bold.
BEIR
Model 135M 500M 1B
Revela 33.6 38.7 42.3
Revela-random 3.9 4.8 4.5
Revela-dual 34.4 38.9 42.9
CoIR
Model 135M 500M 1B
Revela 44.7 52.9 54.6
Revela-random 0.2 0.2 4.8
Revela-dual 43.9 52.2 54.4
Figure 5: Average performance (NDCG@10, %) with and without V-normalization across the retrievers in different sizes. The retrievers with V-normalization perform better than those without the mechanism, signifying the significance of this design.
We investigate the impact of dataset construction strategies, focusing on both batch formation and the composition of the training data.
Batch construction by chunking documents matters.
As previously described, the Revela configuration groups continuous chunks from one or more passages into a single batch, based on the intuition that such chunks are not semantically identical but are contextually related. To assess the importance of this design, we take the training data described in Sec. 3, shuffle all chunks across batches, and regroup them into new, randomly constructed batches. We repeat the same training procedure on both the Wikipedia and code-related corpora to obtain the Revela-random models. As shown in Table 5, the substantial performance drop of Revela-random compared to Revela underscores the importance of structured batch construction. For more detailed results in BEIR and CoIR, please refer to Appendix B.2.
Revela benefits from a larger batch size.
The in-batch attention mechanism in Revela is inspired by the concept of in-batch negatives in supervised contrastive learning xiong2020approximatenearestneighbornegative . As described in Sec. 4.1, the default batch size is set to 16. To analyze the impact of batch size, we construct training data with batch sizes of 8 and 12 using the same batch construction strategy. To ensure a fair comparison, we maintain a consistent total number of training tokens by proportionally increasing the training steps by 100% and 33.3%, respectively. As shown in Fig. 6, Revela achieves the best performance on both BEIR and COIR when trained with a batch size of 16, consistent with the underlying motivation of in-batch negatives. For more detailed results within BEIR and CoIR, please refer to Appendix B.3.
Figure 6: Performance comparison on BEIR and CoIR benchmarks using different batch sizes. For both benchmarks, Revela trained with the batch size of 16 outperforms those trained with 8 and 12.
Revela efficiently learns from mixed-domain composition of the training corpus.
As discussed in Sec. 4.2, when trained separately on Wikipedia and code-related corpora, Revela achieves strong performance on both general and code retrieval benchmarks (BEIR and CoIR, respectively). To explore the potential for scaling up, we examine how combining datasets from mixed domains affects model performance. Specifically, we randomly sample 160,000 batches each from the general and code-related training sets constructed in Sec. 4.1, resulting in a combined set of 320,000 batches. We retrain Revela on the mixed-domain corpus with the same configuration, especially including the number of training steps, and refer to the resulting model as Revela-dual. As shown in Table 5, Revela-dual performs comparably to or even better than the vanilla Revela across different retriever sizes on both BEIR and CoIR, demonstrating the model’s potential for effective generalization across domains.
Figure 7: Comparison of Revela  trained on different recipes.
We further evaluate Revela’s generalization ability by augmenting the Wikipedia data used in Sec. 4.1 with the BEIR corpus. Specifically, we select five domain-specific datasets, including FiQA, NFCorpus, SciFact, SciDocs, and TRECCOVID, and generate a total of 54,312 training batches. These are combined with the existing Wikipedia batches to form a dataset of 320,000 batches. As shown in Fig. 7, models trained on both Wikipedia and BEIR consistently outperform those trained solely on Wikipedia, demonstrating Revela’s effectiveness in adapting to new domains using only the corpus.
5.3 Ablation of V-normalization
In Section 3, we propose a specific design of Revela for in-batch attention computation, which incorporates V-normalization, as defined in Equation 6. The underlying intuition is to prevent the disproportionate amplification of individual patterns within the sequence (e.g., specific tokens or words), thereby encouraging a more comprehensive representation of the overall semantics.
To validate its effectiveness, we conduct an ablation study by jointly training retrievers and LMs without V-normalization. As shown in Figure 5, removing V-normalization leads to a consistent and significant drop in performance across all three Revela variants with different backbone sizes on the BEIR benchmark. These results highlight the critical role of V-normalization in Revela. For a detailed breakdown of per-dataset results, please refer to Appendix B.4.
6 Discussion
Future Directions
This work shows how to train a retriever via self-supervised language modeling on raw text. We outline several future directions to suggest follow-up research. (1) Iterative Indexing: Sec.  5.2 emphasizes the importance of semantically related passages in Revela’s batch construction. While Revela uses document chunking, a more general approach would iteratively index documents and group chunks by on-the-fly representations. Though explored in prior work (izacard2023atlas, ; shi-etal-2024-replug, ), the high computational cost of such methods remains a key challenge for future work. (2) Scaling up: We envision several directions to scale up Revela, including expanding data coverage, as discussed in Sec. 5.2, and increasing model sizes, as analyzed in Sec. 5.1. Additionally, incorporating advanced attention mechanisms (yuan2025native, ) may further enhance retriever learning and accelerate the training process. (3) Potential Influence on LM Training: Revela focuses on retriever learning, treating the LM as auxiliary. LM performance is not analyzed, and it remains unclear whether in-batch attention affects it. Batch construction significantly impacts convergence and model quality (wang2024greats, ; hong2024diversified, ). Future work could explore using retrievers to guide batch construction for LM training.
Limitations
We conclude the limitations of the works from the following perspectives. (1) Evaluation: We only evaluate our methods on an English retrieval benchmark (BEIR) and code retrieval (CoIR). More experiments on multilingual retrieval tasks can make the work more comprehensive. (2) Model scales: As mentioned in Sec. 6, the experiments on larger models (both retrievers and LMs) can better verify the generalization of Revela. (3) Dataset construction: As shown in Sec. 5.2, Revela depends on the intrinsic structure within natural language. How to better leverage the structure for retriever learning is an open question, which has not been covered in this work.
7 Conclusion
Efficiently building information-seeking systems is crucial due to the swiftly evolving world and the wide existence of specific domains. In this work, we introduce Revela, a self-supervised framework that couples dense retrieval with language modeling through a novel in-batch attention mechanism, where a token attends to both local context and other sequences in the batch during NTP. By letting the retriever’s relevance scores weight cross-sequence attention, Revela transforms NTP into a retrieval signal, making use of high-quality and diverse raw text and eliminating the need for manual relevance labels. Our experiments on both general and domain-specific benchmarks demonstrate significant performance gains over existing self-supervised methods, with improvements scaling with retriever size. Further analysis on training data composition and LM size highlights the potential of Revela as a strong and scalable alternative to traditional self-supervised retriever training paradigms, paving the way for more general and efficient retriever learning.
Acknowledgments and Disclosure of Funding
Fengyu Cai is funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science, and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. This work is also supported by the Distr@l4a funding line of the State of Hesse (project number: 493 24_0015_4A). Xinran Zhao is supported by the ONR Award N000142312840.
We thank Kexin Wang, Jan Buchmann, Falko Helm, and Yuhao Zhang for the discussion and comments on an early draft of this work.
Appendix A Experiments
A.1 Evaluation Benchmarks
CoIR contains four classes of code-related retrieval tasks, including Text-to-Code Retrievala, Code-to-Code Retrievalb, Code-to-Text Retrievalc, and Hybrid Code Retrievald. Table 3 presents the specific subtasks included in CoIR with the corresponding descriptions.
Table 3: CoIR Benchmark Tasks. The superscripts present the type of the tasks. The abbreviation of the tasks is noted in the parentheses, presented in Table 2.
Sub-Tasks (Abbr.) Descriptions
AppsRetrieval (Apps)a Retrieve code snippets based on natural language queries.
CosQA (CosQA)a Find code snippets that answer web search queries.
SyntheticText2SQL (ST2SQL)a Retrieve SQL queries based on natural language questions.
COIRCodeSearchNetRetrieval (SN)b Retrieve explanations or summaries for code snippets.
CodeSearchNetCCRetrieval (SNCC)c Identify code snippets that are similar to a given one.
CodeTransOceanContest (TransC)c Retrieve code solutions based on contest problems.
CodeTransOceanDL (TransDL)c Retrieve relevant deep learning code contexts or modules.
StackOverflowQA (SOQA)d Handle hybrid code-related QA tasks with both text and code.
CodeFeedbackST (F-ST)d Retrieve answers to single-turn code-related questions.
CodeFeedbackMT (F-MT)d Handle multi-turn question-answer scenarios in code.
A.2 Examples of the Training Corpus
Fig. 8 presents an example batch containing 16 chunks from two topics. When training, the chunks will be randomly shuffled.
Alcohol – chemistry, properties, and reactions 1. C–C-bond formation and reductions: Barbier, Nozaki–Hiyama, NaBH4/LiAlH4, Meerwein–Ponndorf–Verley, and Noyori asymmetric hydrogenation. 2. Alcohols have pKa 16–19; deprotonation by strong bases yields alkoxides, yet neutral OH is a poor leaving group. 3. Protonation (R–OH $\rightarrow$ R–OH2+) activates SN1/SN2 substitution; e.g. tertiary alcohol + HCl tert-alkyl chloride, or thionyl chloride for primary/secondary cases. 4. Hydrobromic acid or PBr3 give alkyl bromides; Barton–McCombie deoxygenates alcohols … 5. E1 elimination: acid-catalysed dehydration follows Zaitsev, fastest for tertiary alcohols; Fischer esterification and tosylation convert OH into esters. 6. Oxidations: primary aldehyde/carboxylic acid, secondary ketone, tertiary inert; pathway proceeds via hydrate (R–CH(OH)2). 7. Typical oxidants: Collins reagent, Dess–Martin periodinane, KMnO4, Jones reagent; alcohol = any molecule with a C–OH group. Achill Island – geography, history, and culture 8. Largest Irish island (area 148 km2, pop. 2700); linked to the mainland by Michael Davitt Bridge 9. First settlers circa 3000 BC; 87 % peat bog; late-Neolithic population estimate 500–1 000. 10. Deforestation for cultivation; Iron-Age promontory forts at Slievemore, Atlantic Drive, Achillbeg testify to a martial past. 11. Umhall territory ruled by the seafaring O’Malleys; Butler/de Burgo control after Anglo-Norman invasion; 17th–18th-century inward migration. 12. Two Irish dialects co-existed; Carrickkildavnet Castle (15th c.) and the exploits of Grace O’Malley embody maritime heritage. 13. Grace O’Malley met Elizabeth I in 1593; Rev. Edward Nangle founded the proselytising Achill Mission (“the Colony”) in 1831. 14. Mission thrived then waned; Westport–Newport railway …, echoing Brian Rua O’Cearbhain’s prophecy of “carts on iron wheels”. 15. First train carried victims of the Clew Bay drowning (1894); ..neared with the Kirkintilloch bothy fire (1937). 16. Rail line closed weeks after 1937 … St Dymphna’s holy well stand on the south-east coast.
Figure 8: Example of one batch containing chunks split from Wikipedia: This batch contains chunks from two topics: Alcohol (blue) and Achill Island topic (green).
A.3 REPLUG
Self-supervised batch loss.
In our setup, as there are no external queries, each document plays the role of both query and context. For a batch $\mathcal{D}=\{D_{1},\dots,D_{B}\}$, we embed every document with the retriever $E_{\Theta}$:
$s_{ij}\;=\;\frac{E_{\Theta}(D_{i})^{\!\top}E_{\Theta}(D_{j})}{\tau_{r}},\qquad P_{\Theta}(j\mid i,\mathcal{D})=\frac{\exp(s_{ij})}{\sum_{k=1}^{B}\exp(s_{ik})}.$
A frozen LM $g_{\Phi}$ measures how well $D_{j}$ explains $D_{i}$:
$\ell_{ij}\;=\;-\!\log g_{\Phi}(D_{i}\mid D_{j}),\qquad P_{\Phi}(j\mid i,\mathcal{D})=\frac{\exp(-\ell_{ij}/\tau_{\!lm})}{\sum_{k=1}^{B}\exp(-\ell_{ik}/\tau_{\!lm})}.$
We train the retriever by aligning these two distributions for every anchor document:
$\mathcal{L}(\Theta)\;=\;\frac{1}{B}\sum_{i=1}^{B}\operatorname{KL}\!\bigl{(}P_{\Phi}(\cdot\mid i,\mathcal{D})\;\|\;P_{\Theta}(\cdot\mid i,\mathcal{D})\bigr{)}.$
LM parameters $\Phi$ remain fixed; only $\Theta$ is optimized.
Experimental setups
Both of temperatures, $\tau_{\!lm}$ and $\tau_{r}$, are set as 0.001. REPLUG is training on the same datasets with Revela, with the learning rate $5e^{-4}$ and the training steps 4500. All other experimental setups are identical to Revela.
A.4 Checkpoints
We include the off-the-shelf unsupervised and supervised retrievers and the corresponding huggingface URLs in Table 4.
Table 4: Baseline retrievers, LMs (Revela’s backbone), CodeRAG-Bench datasets, and evaluation benchmarks with their HuggingFace URLs and licenses.
Name URL License
Contriever facebook/contriever Not specified
RetroMAE Shitao/RetroMAE Not specified
GraphCodeBERT microsoft/graphcodebert-base Not specified
UniXcoder microsoft/unixcoder-base Apache-2.0
LaPraDoR canwenxu/laprador Apache-2.0
SmolLM2-135M HuggingFaceTB/SmolLM2-135M Apache-2.0
Qwen 2.5-0.5B Qwen/Qwen2.5-0.5B Apache-2.0
LLaMA-3.2-1B meta-llama/Llama-3.2-1B Llama 3.x Community License
Library Documentation code-rag-bench/library-documentation CC BY-SA 4.0
Online Tutorials code-rag-bench/online-tutorials CC BY-SA 4.0
StackOverflow Posts code-rag-bench/stackoverflow-posts CC BY-SA 4.0
BEIR BeIR/beir CC BY-SA 4.0
CoIR CoIR-Retrieval Not specified
A.5 Implementation of Reciprocal Rank Fusion
To combine the strengths of BM25 and Revela, we employ Reciprocal Rank Fusion (RRF) [42]. Given their ranked lists $R_{1}$ and $R_{2}$ for the same query, respectively, RRF assigns a fused score to every document $d$ according to
$\text{score}(d)\;=\;\frac{1}{k+\mathrm{rank}_{R_{1}}(d)}+\frac{1}{k+\mathrm{rank}_{R_{2}}(d)},$
where $\mathrm{rank}_{R_{i}}(d)$ is the $1$-based position of $d$ in run $R_{i}$ (undefined documents are ignored) and $k$ is a constant that controls how quickly the contribution of lower-ranked documents decays; we follow common practice and set $k=60$. After summing the two reciprocal ranks, the documents are re-sorted by this fused score, and the top 200 are retained as the final ranking delivered to RRF.
Appendix B Analysis
B.1 Analysis on the Size of LMs
Tables 5 and 6 present an ablation study analyzing the impact of differently sized LMs.
Table 5: NDCG@10 (%) on BEIR for three model sizes (Revela-135M, Revela-500M, Revela-1B), with LM sizes in ascending order. Each triplet of columns highlights the highest value in bold.
Revela-135M Revela-500M Revela-1B
Dataset 0.1B 0.5B 1B 0.1B 0.5B 1B 0.1B 0.5B 1B
ArguAna 36.5 39.6 39.4 38.0 41.5 41.4 41.0 42.9 45.1
C-FEVER 14.4 16.6 15.6 17.8 18.4 13.5 14.4 15.1 15.6
DBPedia 17.5 17.0 18.3 24.2 22.7 21.4 24.9 26.4 27.6
FEVER 26.3 32.4 35.1 52.9 55.6 50.9 58.2 56.2 60.4
FiQA2018 16.9 19.2 19.1 26.4 24.8 26.6 29.2 27.9 29.6
HotpotQA 41.7 41.1 38.5 54.7 53.9 50.6 58.4 58.7 56.0
NFCorpus 20.0 22.3 23.5 27.3 27.3 26.5 26.3 28.9 27.2
NQ 22.4 21.0 21.2 34.5 32.2 29.8 36.8 34.9 33.9
QuoraRetrieval 79.2 79.8 81.0 83.4 83.5 83.1 82.4 83.3 83.5
SCIDOCS 10.8 11.3 12.1 15.3 15.1 14.8 15.9 16.0 16.4
SciFact 54.1 57.9 57.9 65.5 65.4 65.3 71.6 71.8 71.8
TRECCOVID 58.5 60.9 60.2 68.1 68.8 58.4 66.2 65.4 59.8
Touche2020 11.6 14.9 14.4 27.4 21.5 21.1 26.2 25.0 23.3
Mean 31.5 33.4 33.6 41.2 40.8 38.7 42.4 42.5 42.3
Table 6: Comparison of retrieval performance (%) across CoIR benchmark tasks for three LM model sizes. Bold in each row indicates the best performance in each task for one specific size of retriever.
Revela-135M Revela-500M Revela-1B
Task 0.1B 0.5B 1B 0.1B 0.5B 1B 0.1B 0.5B 1B
AppsRetrieval 5.2 5.9 6.3 15.8 17.9 19.0 12.7 19.6 16.8
CosQA 22.8 22.6 24.2 25.7 25.5 27.6 26.5 26.8 27.7
SyntheticText2SQL 43.2 44.8 44.7 54.4 54.1 53.3 49.7 53.2 53.4
COIRCodeSearchNetRetrieval 36.8 40.5 43.2 47.9 48.1 52.6 53.7 53.7 60.1
CodeSearchNetCCRetrieval 52.2 56.7 58.9 57.6 61.2 63.6 59.6 63.2 67.0
CodeTransOceanContest 60.1 60.1 60.6 68.1 69.0 71.3 67.1 72.0 74.4
CodeTransOceanDL 33.3 33.4 32.9 32.8 33.7 34.1 32.8 32.4 33.7
StackOverflowQA 68.1 66.3 68.3 79.1 77.8 81.0 80.2 81.7 83.8
CodeFeedbackST 63.2 63.0 65.0 72.7 72.5 73.8 72.7 74.6 74.5
CodeFeedbackMT 42.4 42.1 43.1 51.1 51.8 52.3 51.2 54.4 54.6
Mean 42.7 43.5 44.7 50.5 51.2 52.9 50.6 53.2 54.6
From Table 7, we can observe that larger rank in LoRA will lead to better performance for Revela training. This demonstrates the potential scalability of Revela, when we enlarge the rank or conduct full-parameter tuning.
Table 7: Retrieval performance (NDCG@10, %) across BEIR benchmarks, using three model sizes (columns in blocks) and different LoRA ranks (r=16, 64, 256). Bold in each row indicates the highest performance for the specific task and that model size.
Revela-135M Revela-500M Revela-1B
Dataset r=16 r=64 r=256 r=16 r=64 r=256 r=16 r=64 r=256
ArguAna 38.1 38.8 39.4 41.1 40.5 41.4 44.0 44.4 45.1
ClimateFEVER 15.3 15.8 15.6 14.2 13.7 13.5 14.0 13.9 15.6
DBPedia 17.7 17.3 18.3 22.9 19.7 21.4 22.3 22.4 27.6
FEVER 33.1 35.3 35.1 49.2 47.4 50.9 54.6 59.2 60.4
FiQA2018 19.3 18.7 19.1 27.0 25.6 26.6 27.2 28.9 29.6
HotpotQA 38.5 35.6 38.5 51.6 48.0 50.6 52.7 52.6 56.0
NFCorpus 22.9 24.1 23.5 27.8 26.5 26.5 28.1 29.1 27.2
NQ 19.5 19.8 21.2 29.2 25.5 29.8 27.8 28.4 33.9
QuoraRetrieval 81.3 81.9 81.0 83.9 81.5 83.1 82.6 81.8 83.5
SCIDOCS 11.5 11.0 12.1 15.8 15.5 14.8 15.3 16.1 16.4
SciFact 55.9 57.5 57.9 64.4 64.8 65.3 68.2 70.8 71.8
TRECCOVID 57.6 61.2 60.2 61.0 57.4 58.4 59.1 61.9 60.0
Touche2020 14.6 14.1 14.4 20.9 19.2 21.1 18.8 21.5 23.3
Mean 32.7 33.1 33.6 39.2 37.3 38.7 39.6 40.8 42.3
B.2 Analysis on Batch Construction
Table 8 and 9 present the performance of Revela with different setups of training dataset construction on BEIR and CoIR, respectively.
Table 8: Retrieval performance (NDCG@10, %) across BEIR, for three model sizes (135M, 500M, 1B) and different training data setups (vanilla, random, dual).
Revela-135M Revela-500M Revela-1B
Dataset vanilla random dual vanilla random dual vanilla random dual
ArguAna 39.4 13.9 41.6 41.4 10.7 44.4 45.1 15.4 44.7
ClimateFEVER 15.6 6.3 16.3 13.5 7.9 15.5 15.6 6.6 15.1
DBPedia 18.3 1.2 18.7 21.4 2.0 21.7 27.6 3.2 25.4
FEVER 35.1 1.0 36.6 50.9 0.8 44.7 60.4 2.3 56.6
FiQA2018 19.1 0.9 21.0 26.6 2.3 27.9 29.6 1.6 34.0
HotpotQA 38.5 1.1 36.9 50.6 1.4 50.3 56.0 1.4 59.1
NFCorpus 23.5 2.6 24.7 26.5 3.7 28.4 27.2 4.3 28.7
NQ 21.2 0.5 19.7 29.8 0.6 27.4 33.9 2.0 34.5
QuoraRetrieval 80.9 0.5 82.2 83.1 0.4 85.7 83.5 0.5 85.2
SCIDOCS 12.1 1.3 11.6 14.8 2.5 15.7 16.3 2.2 16.1
SciFact 57.9 11.0 58.9 65.2 7.6 62.8 71.8 7.5 69.3
TRECCOVID 60.2 7.6 62.0 58.4 13.2 62.6 59.8 8.2 66.5
Touche2020 14.4 2.8 17.2 21.0 9.3 18.5 23.3 3.6 23.0
Mean 33.6 3.9 34.4 38.7 4.8 38.9 42.3 4.5 42.9
Table 9: Retrieval performance (%) across CoIR, for three model sizes (135M, 500M, 1B) and different training data setups (vanilla, random, dual).
Revela-135M Revela-500M Revela-1B
Dataset chunk random dual chunk random dual chunk random dual
AppsRetrieval 6.3 0.0 6.3 19.0 0.0 17.1 16.8 0.4 17.7
CosQA 24.2 0.0 23.7 27.6 0.0 27.0 27.7 5.2 25.3
SyntheticText2SQL 44.7 0.1 45.9 53.3 0.0 53.4 53.4 1.5 55.2
COIRCodeSearchNetRetrieval 43.2 0.0 41.3 52.6 0.0 51.6 60.1 0.9 58.3
CodeSearchNetCCRetrieval 58.9 0.0 59.3 63.6 0.0 62.7 66.9 1.2 66.6
CodeTransOceanContest 60.6 1.0 57.3 71.3 0.5 70.7 74.4 3.8 74.1
CodeTransOceanDL 32.9 0.7 34.2 34.1 1.2 33.8 33.7 8.3 34.0
StackOverflowQA 68.3 0.3 66.8 81.0 0.0 80.0 83.8 19.7 83.3
CodeFeedbackST 65.0 0.0 63.0 73.8 0.0 73.5 74.5 4.4 74.5
CodeFeedbackMT 43.1 0.1 41.9 52.3 0.0 52.0 54.6 2.7 55.3
Mean 44.7 0.2 44.0 52.9 0.2 52.2 54.6 4.8 54.4
B.3 Analysis on Batch sizes
Table 10 and 11 present the performance of Revela with different batch sizes on BEIR and CoIR, respectively. Although models trained on a batch size of 16 achieve higher mean performance, they cannot consistently outperform those trained with smaller batch sizes across all tasks. This finding demonstrates the possibility to adopt smaller batch sizes for Revela training.
Table 10: Retrieval performance (NDCG@10, %) across the 13 BEIR tasks for three encoder sizes (135M, 500M, 1B) and three batch sizes (bs8, bs12, bs16).
Revela-135M Revela-500M Revela-1B
Dataset bs8 bs12 bs16 bs8 bs12 bs16 bs8 bs12 bs16
ArguAna 39.7 41.1 39.4 42.9 39.9 41.4 44.4 44.4 45.1
ClimateFEVER 15.7 15.9 15.6 15.2 12.5 13.5 16.0 13.3 15.6
DBPedia 17.1 17.0 18.3 21.5 18.1 21.4 25.9 23.7 27.6
FEVER 36.0 37.0 35.1 54.2 44.4 50.9 55.4 58.7 60.4
FiQA2018 20.1 18.7 19.1 26.2 25.5 26.6 29.7 29.0 29.6
HotpotQA 36.8 35.6 38.5 50.9 46.3 50.6 55.4 53.9 56.0
NFCorpus 22.7 24.0 23.5 24.1 25.7 26.5 30.1 29.6 27.2
NQ 19.4 19.5 21.2 24.7 22.4 29.8 30.8 30.5 33.9
QuoraRetrieval 80.0 80.2 81.0 82.6 82.2 83.1 82.1 83.4 83.5
SCIDOCS 11.5 11.3 12.1 14.8 14.3 14.8 16.5 15.7 16.3
SciFact 56.9 58.0 57.9 63.2 62.2 65.2 68.8 69.3 71.8
TRECCOVID 61.0 59.7 60.2 57.3 58.5 58.4 56.7 59.8 59.8
Touche2020 16.3 14.1 14.4 20.0 20.7 21.0 24.9 22.9 23.3
Mean 33.3 33.2 33.6 38.3 36.4 38.7 41.3 41.1 42.3
Table 11: Retrieval performance (%, NDCG@10) across the 10 CoIR tasks. We report three encoder sizes (135M, 500M, 1B) and three batch sizes (bs8, bs12, bs16).
Revela-135M Revela-500M Revela-1B
Dataset bs8 bs12 bs16 bs8 bs12 bs16 bs8 bs12 bs16
AppsRetrieval 6.8 6.5 6.3 17.8 16.5 19.0 17.9 19.0 16.8
CosQA 22.2 24.0 24.2 26.1 28.9 27.6 27.9 28.6 27.7
SyntheticText2SQL 42.0 41.9 44.7 52.8 55.2 53.3 54.8 55.5 53.4
COIRCodeSearchNetRetrieval 43.5 43.4 43.2 52.5 50.8 52.6 58.8 56.2 60.1
CodeSearchNetCCRetrieval 57.9 59.8 58.9 62.0 63.3 63.6 66.2 65.5 67.0
CodeTransOceanContest 58.8 58.7 60.6 66.8 70.1 71.3 71.8 74.2 74.4
CodeTransOceanDL 34.7 32.6 32.9 33.5 33.9 34.1 32.9 33.3 33.7
StackOverflowQA 67.4 66.7 68.3 78.2 80.9 81.0 84.8 85.0 83.8
CodeFeedbackST 62.5 63.0 65.0 73.3 74.7 73.8 75.0 75.1 74.5
CodeFeedbackMT 41.6 41.7 43.1 51.8 51.6 52.3 55.1 54.1 54.6
Mean 43.7 43.8 44.7 51.5 52.6 52.9 54.5 54.7 54.6
B.4 Analysis of V-normalization
Table 12 compares Revela with and without the proposed V-normalization. We observe that across various datasets within BEIR, including V-normalization mostly outperforms its absence, highlighting the effectiveness of the design.
Table 12: NDCG@10 (%) on BEIR across three models in different sizes. Bold numbers in each row indicate the better performance between w/o v-norm vs. w/ v-norm for a given model size.
Revela-135M Revela-500M Revela-1B
Dataset w/o v-norm w/ v-norm w/o v-norm w/ v-norm w/o v-norm w/ v-norm
ArguAna 42.5 39.4 41.5 41.4 45.6 45.1
C-FEVER 14.5 15.6 12.0 13.5 12.7 15.6
DBPedia 16.1 18.3 17.1 21.4 22.8 27.6
FEVER 29.0 35.1 41.3 50.9 46.0 60.4
FiQA2018 18.2 19.1 25.1 26.6 26.9 29.6
HotpotQA 35.8 38.5 44.8 50.6 49.7 56.0
NFCorpus 23.7 23.5 26.8 26.5 29.1 27.2
NQ 16.2 21.2 19.1 29.8 22.1 33.9
QuoraRetrieval 81.5 80.9 80.4 83.1 83.3 83.5
SCIDOCS 10.7 12.1 14.6 14.8 16.5 16.3
SciFact 58.2 57.9 62.6 65.2 71.6 71.8
TRECCOVID 47.6 60.2 50.4 58.4 48.1 59.8
Touche2020 11.1 14.4 14.7 21.0 17.6 23.3
Mean 31.2 33.6 34.7 38.7 37.9 42.3